# supplementary

### v1

@@ Abstract
Scaling up deliberative and voting participation is a longstanding endeavor – a cornerstone for direct democracy and legitimate collective choice. Recent breakthroughs in generative artificial intelligence (AI) and large language models (LLMs) provide unprecedented opportunities, but also alerting risks for digital democracy. AI personal assistants can overcome cognitive bandwidth limitations of humans, providing decision support capabilities or even direct AI representation of human voters at large scale. However, the quality of this representation and what underlying biases manifest when delegating collective decision making to LLMs is an alarming and timely challenge to tackle. By rigorously emulating with high realism more than >50K LLM voting personas in 81 real-world voting elections, we show that different LLMs (GPT 3, GPT 3.5, and Llama2) come with biases and significant inconsistencies in complex preferential ballot formats, compared to simpler and more consistent majoritarian elections. Strikingly, fair voting aggregation methods, such as equal shares, prove to be a win-win: fairer voting outcomes for humans with fairer AI representation. This novel underlying relationship proves paramount for democratic resilience in progressives scenarios with low voters turnout and voter fatigue supported by AI representatives: abstained voters are mitigated by recovering highly representative voting outcomes that are fairer. These insights provide remarkable foundations for science, policymakers and citizens in explaining and mitigating AI risks in democratic innovations.
@@ 1 Introduction
Generative AI voting: missing opportunities with inevitable challenges. Large language models have been used to predict election winners using sensitive demographic data about the political profile of individuals [18]. They have also been used to predict pairwise comparisons of proposals for constitutional changes [6] and summarized opinions expressed with free-form text [19]. However, little is known whether this AI predictive capability can expand to more complex collective decision-making processes that involve more options to choose from, with more complex preference elicitation methods [5].
Participatory budgeting [20] is one such process put under scrutiny in this article. Here city authorities distribute a public budget by letting citizens propose their own project ideas, which they vote for and often implement themselves [21]. Projects may be pertinent to different impact areas (e.g. environment, culture, welfare), beneficiaries (e.g. elderly, children) and can have different costs. Voters can approve, rank or distribute points over their preferred projects, while winners are elected based on the popularity of the projects (utilitarian greedy) or based on a proportional representation of the voters’ preferences (equal shares) [22, 23]. Recently though, a reinforcement learning approach has been used to elect winners by assisting voters to reach a consensus [24]. So far, the scope and citizens’ engagement in participatory budgeting campaigns remains highly local [25, 11]. With such complexity and degree of design freedom, scaling up participatory budgeting turns into the ultimate democratic blueprint to assess capabilities and risks of generative AI voting.
@@
How resilient representative voting outcomes are with generative AI. We hypothesize that generative AI representatives of human voters build up resilience for representative voting outcomes in voting processes of low turnout if these human voters would otherwise abstain or would not have capacity to actively participate in multiple collective decisions at large-scale (see Figure 1d. In other words, we explore whether the inconsistencies of voting outcomes because of low voting turnout are higher than inconsistencies originated from generative AI voters.
@@
The inconsistencies of generative AI, along with the potential biases explaining these inconsistencies, are systematically studied here for the first time within a novel factorial design that consists of 5 dimensions (see Figure 1): (i) Real-world voting scenarios - datasets from the 2012, 2016 and 2020 US national elections [31] as well as the mock and actual participatory budgeting voting of the ‘City Idea’ campaign of 2023 in Aarau, Switzerland [32] are studied. They come with different number of ballot options and voters, see Figure 1a and Section 4.2. (ii) Personal human traits - for each voter, multiple incremental levels of additional information are given as input to large language models to generate voting choices. This information includes socio-demographics, political preferences, preferences for project categories and preferences for voting outcome qualities (popular vs. fair election of winners, individual vs. collective benefit). These are derived from survey data collected along the voting campaigns (see Section 4.1, and Table LABEL:table:prologue1 and LABEL:table:prologue2). (iii) Voting input methods - four methods with incremental level of complexity and expressiveness are compared [28, 33]. These include single choice for all voting scenarios, approvals, score (assigning a preference score to each option) and cumulative voting (distributing a number of points over the options) [34, 35, 36] for the participatory budgeting scenarios. (iv) AI models - Three large language models are assessed along with a more mainstream predictive machine learning (ML) model used as a benchmark. GPT 3, GPT 3.5, and Llama2 are chosen covering a large spectrum of capabilities in open and closed generative AI [37], see Section 4.1 for more information. (v) Voting aggregation methods - majority is used for the scenario of US elections. The standard utilitarian greedy and the method of equal shares [38] are used for the participatory budgeting scenarios. The former method iterates to elect the next most popular project as long as the budget is not exhausted. The latter method elects the projects that make the most people ‘happy’, meaning a combination of projects that maximizes a proportional representational of voters’ preferences. This is the method that was actually used in City Idea campaign to elect the winners [39]. See Section 4.1 for more information.
@@
The dimensions of the factorial design are illustrated in Table 1 and the studied combinations are marked with the color boxes in Figure 1a. This broad spectrum of analysis allows us to generalize the findings of the study and make them highly relevant for a broad spectrum of research communities and policymakers.
@@
Figure 1: An overview of the generative AI framework, along with the factorial design and the biases studied in this article. (A) The factorial design with the 5 studied dimensions: (i) real-world voting scenarios in the context of participatory budgeting and national elections. (ii) Various combinations of persona human traits based on which AI personas are created. (iii) Four voting input methods. (iv) Four AI models, three large language models and a predictive machine learning model (benchmark). (v) Several voting aggregation methods for elections and participatory budgeting. The studied combinations for each voting scenario are marked with different colors, see also Table 1. (B) The framework of generative AI voting. For each voter in the real-world voting scenario, a prompt is given to large language models to construct the voting persona. The input is the personal human traits, the voting options and the ballot format with instructions for the voting persona about how to make a choice. This choice is the output of the persona. Both human and AI choices are aggregated using a voting aggregation method. The inconsistencies of individual and collective choices for humans and AI personas are assessed, along with potential biases that explain these inconsistencies. (C) The personal human traits are mapped to cognitive biases. Section 4.4 illustrates choice inconsistencies and their origin to potential cognitive biases. (D) Hypothesizing for a recovery of consistency in voting outcomes under low voters turnout if AI representatives support abstained human voters.
@@
Assessing generative AI voting in action. Voting personas are created using input prompts of large language models as depicted in Figure 1b. This designed process aims to emulate the three voting scenarios with the different settings of Figure 1a. Each input prompt consists of a standardized description of the voter’s profile (see Section LABEL:sec:study in SI) and an instruction to vote according to the ballot format. The individual and collective choices of humans and AI personas are compared by measuring the overlap (Jaccard similarity) of their selected/winning options as an indicator of consistency [28]. These consistency values are then becoming the dependent variable to predict using the personal human traits as independent variables (features), which are fed in a neural network (see Section 4.5). Based on a systematic mapping of human personal traits to the cognitive biases as illustrated in Figure 1c (see Section 4for more detail), this prediction model causally explains the human traits that contribute to inconsistencies and the potential underlying biases that explain these inconsistencies.
@@ 2 Results
1. Fair voting methods to elect winners are more resilient to human-AI inconsistencies in collective choice, demonstrating a novel underlying win-win relationship: fairer voting outcomes for humans with fairer human representation by AI (Figure 2). These inconsistencies are particularly prominent in complex ballot formats with a large number of alternatives, while simple majoritarian voting tends to be highly consistent. GPT 3.5 shows transitivity of preferences among different ballot formats (Figure 3).
@@
3. Conformity biases explain human-AI choice consistency, while unconscious biases explain inconsistency (Figure 5). Political/ideological profile determines the human-AI consistency in single-choice majoritarian voting with two candidates. Time discounting factors explain transitivity of AI preferences between different ballot formats, while affect heuristics and unconscious biases explain intransitivity.
@@ 2.1 Fair collective choice is resilient to human-AI inconsistencies
Figure 2 illustrates the human-AI consistency in individual and collective choices. The consistency of individual choice remains poor in complex ballot formats with several alternatives. On average, it is 4.5% and 27.2% for the actual and mock voting in participatory budgeting, yet it is 82.3% for the binary majoritarian US elections. GPT 3.5 shows the highest consistency of individual choice among the large language models, which is 5.2% and 5.31% higher than GPT 3 and Llama2. On the contrary, the consistency of the collective choice increases by 51.2% in total. Strikingly, the consistency of equal shares is 80.8%, which is 32.8% higher than utilitarian greedy. Even when reducing the number winners of equal shares to the 8 of utilitarian greedy, the consistency remains 29.4% higher.
Figure 2: Choice by large language models is consistent to humans for single-choice majoritarian elections, however consistency drops for more complex ballots with larger number of alternatives as in the case of participatory budgeting. Strikingly, consistency of collective choice is significantly higher than individual choice, particularly for the fairer voting aggregation rule of equal shares. GPT 3.5 shows the highest choice consistency and Llama2 the lowest among the large language models, which though remain inferior to a predictive machine learning model. The consistency (y-axis) in individual and collective choice is shown for different AI models (x-axis) across three real-world voting scenarios: The participatory budgeting campaign of City Idea, (A) actual and (B) mock, and (C) the US national elections of 2012, 2016 and 2020. The consistency in the participatory scenarios is shown for different voting input methods (left and right). The consistency of collective choice with the two voting aggregation methods of utilitarian greedy and equal shares is also shown. For the actual voting of City Idea, the consistency of equal shares is calculated for all winners and 8 winners (as many as greedy) for a fair comparison.
Compared to the large language models, the machine learning model shows 3.6% higher consistency in individual choice and 6.6% higher in collective choice. All consistency values are the maximum ones derived with all personal human traits as shown in Figure 1a. The removal of project preferences shows the highest consistency reduction of 19.7%, while political interest shows the lowest consistency reduction of 3.4%.
Figure 3 illustrates the transitivity of preferences by different large language models by measuring the consistency of individual choice among different pairs of ballot formats. ‘Approval - cumulative’ voting shows the highest consistency, which is 1.8% higher than ‘single choice - approval’ and 3.5% higher than ‘single choice - cumulative’.
@@ 2.2 AI representatives to recover from low voters turnout
Figure 4a illustrates the capability in AI representatives who abstain from voting to recover the consistency of the voting outcomes lost by low voters turnouts. Strikingly, up to 50% of AI representatives are sufficient for a maximum consistency recovery of 53.3% under voting turnouts lower than 80% when equal shares is used. This significant recovery is despite the very low consistency of human-AI individual choices in complex ballot formats with several options (Figure 2a). The recovery is also confirmed for equal shares with the top-8 winners, which is equal to the number of winners with utilitarian greedy (on average 18.1% lower consistency recovery). However, AI representatives fail to recover consistency with utilitarian greedy as the voting aggregation method.
Figure 4: By representing up to half of the human voters that abstained from voting with AI, a significant lost consistency in the collective choice is recovered in participatory budgeting under fair voting aggregation methods such as equal shares and under low number of alternatives. However, further consistency loss is observed under a large number of alternatives with the standard utilitarian greedy method as the number of AI representatives increases. In contrast, single-choice majoritarian elections with highly consistent human-AI individual choice maximize consistency recovery for the whole spectrum of voters turnout by maximizing the AI representatives for the voters who abstain. Nevertheless, for more than 50% of voters turnout, consistency is fully recovered by representing at least 25% of abstained voters with AI. The consistency loss in the voting outcome by low voters turnout (x-axis) is emulated by removing different ratios of human voters. A recovery of consistency (y-axis) is hypothesized by different ratios of abstained voters represented by AI: 25%, 50%, 75% and 100%. The consistency recovery values are the means of all three large language models (see Figure LABEL:fig:LLM_factorial for the results of each large language model). The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea are shown with two voting aggregation methods: utilitarian greedy and equal shares, with all winners and the top-8 (as many as greedy) for a fairer comparison. (C) US elections. (D) Chord visualizations summarizing how the interplay between voters turnout and AI representation yields voting outcomes with consistency gain or loss.
@@ 2.3 Biases explaining AI inconsistencies in choice and preference transitivity
To explain the origin of (in)consistencies in human-AI choice and different voting input methods, a prediction model is designed for each voting scenario. The personal human traits, mapped to different cognitive biases, become the independent variables and the consistency the dependent variable. The prediction model is built using a neural network (see Section 4.4). The relative importance of the personal human traits that explain the AI consistency is calculated using category gradient boosting (see Section 4.5).
@@
In the case of participatory budgeting in Figure 5a and 5b, the preference for environmental projects (potential conformity bias [40, 41, 42]) shows an average importance of 26.5% (p=0.0258) in explaining human-AI choice consistency (25% and 12% higher consistency among environmental supporters, with p=0.038 for the actual voting). Political belief and ideology are strongly correlated with the actual two alternatives in the US elections (right/conservative for republicans and left/liberal for democrats), therefore, they strongly explain the consistency of human-AI choice (mean relative importance of 98.7%, p<0.02 and 4.8%, p<0.03 respectively). However, this is not the case in participatory budgeting as it political belief that rather explains inconsistency (relative importance of -51.2% with p<0.02). This demonstrates that the underlying participatory budgeting choices remain politically neutral to a high extent, despite the fact that the City Idea campaign mainly attracted left-wing voters (42.5%) versus 31.5% of right wing voters. In contrast, race and discussion of politics explain inconsistencies in the US elections with a mean relative importance of -48.1% (p<0.04) and -64.7% (p<0.04) respectively. Inconsistency is 33% higher than consistency among white people (p<0.019), while for people who discuss politics inconsistency is 63.2% higher that the ones who do not (p=0.046).
@@
Figure 5: Conformity bias, such as preference for environmental projects, positively contribute to the consistency of human-AI choice. On the other hand, unconscious biases such as race, discussion of politics and political belief cause inconsistencies. In the case of the US elections though, political belief and ideology strongly contribute to consistency. Time discounting factors such as preference for sport, public space and welfare projects positively contribute to the consistency of large language models among the voting input methods. However, affect heuristics such as preference for projects that benefit elderly cause inconsistencies among voting methods. Unconscious biases originated from socio-demographic information also cause inconsistencies in the simple mock voting scenario. The relative importance (x-axis) of the personal human traits (y-axis) to predict the consistency of choice and consistency between voting input methods (single choice vs. cumulative/score) is shown for different large language models and the machine learning model. This is calculated using category gradient boosting. The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea is shown along with the (C) US elections. The ‘*’ denotes that the feature is statistically significant (p<0.05) for each of the AI models of Llama2, GPT 3, GPT 3.5 and ML respectively. The chord visualizations summarize the biases that explain the (in)consistencies of (D) human-AI choice and (E) voting input methods.
@@
The consistency among the voting input methods is explained by preference towards time discounting factors [41] in participatory budgeting, such as sport, welfare and public space projects. The average relative importance is 25.6%, 25.3% and 17.5% respectively (all with p<0.03), with 48.53%, 44.64%, and 37.5% of higher consistency among supporters in the actual voting (p<0.044). However, support for elderly, linked to potential affect heuristics [41, 43, 44], explains inconsistency with an average relative importance of -35.4% (56.4% higher inconsistency among elderly supporters, with p=0.0371 for the actual voting).
Figures LABEL:fig:explaination-consistency1LABEL:fig:merged3aLABEL:fig:explaination-top-choice, and Table LABEL:table:error illustrate additional insights about how personal human traits explain the AI top choice and the human-AI consistency of the individual choices.

### v2

@@ Abstract
Scaling up deliberative and voting participation is a longstanding endeavor – a cornerstone for direct democracy and legitimate collective choice. Recent breakthroughs in generative artificial intelligence (AI) and large language models (LLMs) unravel new capabilities for AI personal assistants to overcome cognitive bandwidth limitations of humans, providing decision support or even direct representation of human voters at large scale. However, the quality of this representation and what underlying biases manifest when delegating collective decision making to LLMs is an alarming and timely challenge to tackle. By rigorously emulating with high realism more than >50K LLM voting personas in 81 real-world voting elections, we disentangle the nature of different biases in LLMS (GPT 3, GPT 3.5, and Llama2). Complex preferential ballot formats exhibit significant inconsistencies compared to simpler majoritarian elections that show higher consistency. Strikingly though, by demonstrating for the first time in real-world a proportional representation of voters in direct democracy, we are also able to show that fair ballot aggregation methods, such as equal shares, prove to be a win-win: fairer voting outcomes for humans with fairer AI representation. This novel underlying relationship proves paramount for democratic resilience in progressives scenarios with low voters turnout and voter fatigue supported by AI representatives: abstained voters are mitigated by recovering highly representative voting outcomes that are fairer. These interdisciplinary insights provide remarkable foundations for science, policymakers and citizens to develop safeguards and resilience for AI risks in democratic innovations.
@@ 1 Introduction
Generative AI voting: a converging advance with inevitable challenges. Large language models have been used to predict election winners using sensitive demographic data about the political profile of individuals [18]. They have also been used to predict pairwise comparisons of proposals for constitutional changes [6] and summarized opinions expressed with free-form text [19]. However, little is known whether this AI predictive capability can expand to more complex collective decision-making processes that involve more options to choose from, with more complex preference elicitation methods [5].
Participatory budgeting [20] is one such process put under scrutiny in this article. Here city authorities distribute a public budget by letting citizens propose their own project ideas, which they vote for and often implement themselves [21]. Projects may be pertinent to different impact areas (e.g. environment, culture, welfare), beneficiaries (e.g. elderly, children) and can have different costs. Voters can approve, rank or distribute points over their preferred projects, while winners are elected based on the popularity of the projects (utilitarian greedy) or based on a proportional representation of the voters’ preferences (equal shares[22, 23]. Recently though, a reinforcement learning approach has been used to elect winners by assisting voters to reach a consensus [24]. So far, the scope and citizens’ engagement in participatory budgeting campaigns remains to a large extent a one-shot and highly local [25, 11]. What if there was a way to geographically spread such decision-making processes and run them on a regular basis? With such complexity and degree of design freedom, scaling up participatory budgeting turns into the ultimate democratic blueprint to assess capabilities and risks of generative AI voting.
@@
How resilient representative voting outcomes are with generative AI. We hypothesize that generative AI representatives of human voters build up resilience for representative voting outcomes in voting processes of low turnout if these human voters would otherwise abstain or would not have capacity to actively participate (see Figure 1d). In other words, we explore whether the inconsistencies of voting outcomes because of low voting turnout are higher than inconsistencies originated from generative AI representatives of abstained voters.
@@
The inconsistencies of generative AI, along with the potential biases explaining these inconsistencies, are systematically studied here for the first time using a novel factorial design based on real-world empirical evidence. It consists of 5 dimensions (see Figure 1): (i) Real-world voting scenarios - datasets from the 2012, 2016 and 2020 US national elections [31] as well as the mock and actual participatory budgeting voting of the ‘City Idea’ campaign of 2023 in Aarau, Switzerland [32] are studied. They come with different number of ballot options and voters, see Figure 1a and Section 4.2. (ii) Personal human traits - for each voter, multiple incremental levels of additional information are given as input to large language models to generate ballots. This information includes socio-demographics, political preferences, preferences for project categories and preferences for qualities of voting outcomes (popular vs. fair election of winners, individual vs. collective benefit). These are derived from real-world evidence including survey data collected along the voting campaigns (see Section 4.1, and Table LABEL:table:prologue1 and LABEL:table:prologue2). (iii) Voting input methods - four methods with incremental level of complexity and expressiveness are compared [28, 33]. These include single choice for all voting scenarios, approvals, score (assigning a preference score to each option) and cumulative voting (distributing a number of points over the options) [34, 35, 36] for the participatory budgeting scenarios. (iv) AI models - Three large language models are assessed along with a more mainstream predictive machine learning (ML) model used as a benchmark. GPT 3, GPT 3.5, and Llama2 are chosen covering a large spectrum of capabilities in open and closed generative AI [37], see Section 4.1 for more information. (v) Voting aggregation methods - majority is used for the scenario of US elections. The standard utilitarian greedy and the method of equal shares [38] are used for the participatory budgeting scenarios. The former method iterates to elect the next most popular project as long as the budget is not exhausted. It has been the standard method used in real-world. The latter method elects the projects that make the most people ‘happy’, meaning a combination of projects that maximizes a proportional representational of voters’ preferences. This is the method that was actually used in City Idea campaign to elect the winners [39], providing significance and realism for the findings of this study. See section 4.1 for more information.
@@
The dimensions of the factorial design are illustrated in Table 1 and the studied combinations are marked with the colored boxes in Figure 1a. This broad spectrum of analysis based on real-world evidence allows us to generalize the findings of the study and make them highly relevant for a broad spectrum of research communities and policymakers.
@@
Figure 1: An overview of the generative AI framework, along with the factorial design and the biases studied in this article. (A) The factorial design with the 5 studied dimensions: (i) real-world voting scenarios in the context of participatory budgeting and national elections. (ii) Various combinations of persona human traits based on which AI personas are created. (iii) Four voting input methods. (iv) Four AI models, three large language models and a predictive machine learning model (benchmark). (v) Voting aggregation methods for elections and participatory budgeting. The studied combinations for each voting scenario are marked with different colors, see also Table 1. (B) The framework of generative AI voting. For each voter in the real-world voting scenario, a prompt is given to large language models to construct the voting persona. The input is the personal human traits, the voting options and the ballot format with instructions for the voting persona about how to make a choice. This choice is the output of the persona. Both human and AI choices are aggregated using a voting aggregation method. The inconsistencies of individual and collective choices for humans and AI personas are assessed, along with potential biases that explain these inconsistencies. (C) The personal human traits are mapped to cognitive biases. Section 4.4 illustrates choice inconsistencies and their origin to potential cognitive biases. (D) Hypothesizing for a recovery of consistency in voting outcomes under low voters turnout using AI representatives for abstained human voters.
@@
Assessing generative AI voting in action. Voting personas are constructed using input prompts of large language models as depicted in Figure 1b. This designed process aims to emulate the three voting scenarios with the different settings of Figure 1a. Each input prompt consists of a standardized description of the voter’s profile (see Section LABEL:sec:study in SI) and an instruction to vote according to the ballot format. The individual and collective real-world choices of humans and AI personas are compared for the first time by measuring the overlap (Jaccard similarity) of their selected/winning options as an indicator of consistency [28]. These consistency values are then becoming the dependent variable to predict using the personal human traits as independent variables (features), which are fed in a neural network (see Section 4.5). Based on a systematic mapping of human personal traits to the cognitive biases as illustrated in Figure 1c (see Section 4 for more detail), this prediction model causally explains the human traits that contribute to inconsistencies and the potential underlying biases that explain these inconsistencies. This novel analysis is designed to provide a significant conceptual advance to understand how voting design reinforces or mitigates different AI biases in real-world practice.
@@ 2 Results
1. Fair voting methods to elect winners are more resilient to human-AI inconsistencies, demonstrating a striking underlying win-win relationship: fairer voting outcomes for humans with fairer human representation by AI (Figure 2). These inconsistencies are particularly prominent in complex ballot formats with a large number of alternatives, while simple majoritarian voting tends to be highly consistent. GPT 3.5 shows transitivity of preferences among different ballot formats (Figure 3).
@@
3. Conformity biases explain consistency of human-AI choice, while unconscious biases explain inconsistency (Figure 5). Political/ideological profile determines the human-AI consistency in single-choice majoritarian voting with two candidates. Time discounting factors explain transitivity of AI preferences between different ballot formats, while unconscious biases explain intransitivity.
@@ 2.1 Fair collective choice is resilient to human-AI inconsistencies
Figure 2 illustrates the human-AI consistency in individual and collective choices. The consistency of individual choice remains poor in complex ballot formats with several alternatives. On average, it is 4.5% and 27.2% for the actual and mock voting in participatory budgeting, yet it is 82.3% for the binary majoritarian US elections. GPT 3.5 shows the highest consistency of individual choice among the large language models, which is 5.2% and 5.31% higher than GPT 3 and Llama2. On the contrary, the consistency of the collective choice increases by 51.2% in total. Strikingly, the consistency of equal shares is on average 80.8%, which is 32.8% higher than utilitarian greedy. Even when reducing the number winners of equal shares to the 8 of utilitarian greedy, the consistency remains 29.4% higher.
Figure 2: Choice by large language models is consistent to humans for single-choice majoritarian elections, however consistency drops for more complex ballots with larger number of alternatives as in the case of participatory budgeting. Strikingly, consistency of collective choice is significantly higher than individual choice, particularly for the fairer voting aggregation rule of equal shares. GPT 3.5 shows the highest choice consistency and Llama2 the lowest among the large language models, which though remain inferior to a predictive machine learning model. The consistency (y-axis) in individual and collective choice is shown for different AI models (x-axis) across three real-world voting scenarios: The participatory budgeting campaign of City Idea, (A) actual and (B) mock, and (C) the US national elections of 2012, 2016 and 2020. The consistency in the participatory budgeting scenarios is shown for different voting input methods (left and right). The consistency of collective choice with the two voting aggregation methods of utilitarian greedy and equal shares is also shown. For the actual voting of City Idea, the consistency of equal shares is calculated for all winners and 8 winners (as many as greedy) for a fairer comparison.
Compared to large language models, the machine learning model shows 3.6% higher consistency in individual choice and 6.6% higher in collective choice. All consistency values are the maximum ones derived with all personal human traits as shown in Figure 1a. The removal of project preferences shows the highest consistency reduction of 19.7%, while political interest shows the lowest consistency reduction of 3.4%.
Figure 3 illustrates the transitivity of preferences in different large language models by measuring the consistency of individual choice among different pairs of ballot formats. ‘Approval - cumulative’ voting shows the highest consistency, which is 1.8% higher than ‘single choice - approval’ and 3.5% higher than ‘single choice - cumulative’.
@@ 2.2 AI representatives to recover from low voters turnout
Figure 4a illustrates the capability of AI representatives to recover the consistency of voting outcomes lost by low voters turnouts. Strikingly, up to 50% of AI representatives are sufficient for a maximum consistency recovery of 53.3% under voting turnouts lower than 80% when equal shares is used. This significant recovery is despite the very low consistency of human-AI individual choices in complex ballot formats with several options (Figure 2a). The recovery is also confirmed for equal shares with the top-8 winners, which is equal to the number of winners with utilitarian greedy (on average 18.1% lower consistency recovery). However, AI representatives fail to recover consistency with utilitarian greedy as the voting aggregation method.
Figure 4: By representing up to half of the human voters who abstain from voting with AI, a significant lost consistency in the collective choice is recovered in participatory budgeting under fair voting aggregation methods such as equal shares and under low number of alternatives. However, further consistency loss is observed under a large number of alternatives with the standard utilitarian greedy method as the number of AI representatives increases. In contrast, single-choice majoritarian elections with highly consistent human-AI individual choice maximize consistency recovery for the whole spectrum of voters turnout by maximizing the AI representatives for the voters who abstain. Nevertheless, for more than 50% of voters turnout, consistency is fully recovered by representing at least 25% of abstained voters with AI. The consistency loss in the voting outcome by low voters turnout (x-axis) is emulated by removing different ratios of human voters. A recovery of consistency (y-axis) is hypothesized by different ratios of abstained voters represented by AI: 25%, 50%, 75% and 100%. The consistency recovery values are the means of all three large language models (see Figure LABEL:fig:LLM_factorial for the results of each large language model). The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea are shown with two voting aggregation methods: utilitarian greedy and equal shares, with all winners and the top-8 (as many as greedy) for a fairer comparison. (C) US elections. (D) Chord visualizations summarizing how the interplay between voters turnout and AI representation yields voting outcomes with consistency gain or loss. (E) Visualizations summarizing the mechanism of consistency recovery through AI representatives for the voters who abstain.
@@ 2.3 Biases explaining AI inconsistencies in choice and preference transitivity
To explain the origin of (in)consistencies in human-AI choice and different voting input methods, a prediction model is designed for each voting scenario. The personal human traits, mapped to different cognitive biases, become the independent variables and the consistency the dependent variable. The prediction model is built using a neural network (see Section 4.4). The relative importance of the personal human traits that explain the AI consistency is calculated using shapley additive explanations (see Section 4.5).
@@
In the case of participatory budgeting in Figure 5a and 5b, the preference for environmental projects (potential conformity bias [40, 41, 42]) shows an average importance of 26.5% (p=0.0258) in explaining human-AI choice consistency (25% and 12% higher consistency among environmental supporters, with p=0.038 for the actual voting scenario). Political belief and ideology are strongly correlated with the actual two alternatives in the US elections (right/conservative for republicans and left/liberal for democrats), therefore, they strongly explain the consistency of human-AI choice (mean relative importance of 98.7%, p<0.02 and 4.8%, p<0.03 respectively). However, this is not the case in participatory budgeting as it is political belief that rather explains inconsistency (relative importance of -51.2% with p<0.02). This demonstrates that the underlying participatory budgeting choices remain politically neutral to a high extent, despite the fact that the City Idea campaign mainly attracted left-wing voters (42.5% versus 31.5% of right wing voters). In contrast, race and discussion of politics explain inconsistencies in the US elections with a mean relative importance of -48.1% (p<0.04) and -64.7% (p<0.04) respectively. Inconsistency is 33% higher than consistency among white people (p<0.019), while for people who discuss politics inconsistency is 63.2% higher that the ones who do not (p=0.046).
@@
Figure 5: Conformity bias, such as preference for environmental projects, positively contribute to the consistency of human-AI choice. On the other hand, unconscious biases such as race and discussion of politics cause inconsistencies. For US elections though, political belief and ideology strongly contribute to consistency. Time discounting factors such as preference for sport, and welfare projects as well as affect heuristics such as preference for projects that benefit families positively contribute to the consistency of large language models among the voting input methods. The relative importance (x-axis) of the personal human traits (y-axis) to predict the consistency of choice and consistency between voting input methods (single choice vs. cumulative/score) is shown for different large language models and the machine learning model. This is calculated using Shapley Additive Explanations. The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea is shown along with the (C) US elections. The ‘*’ denotes that the feature is statistically significant (p<0.05) for each of the AI models of Llama2, GPT 3, GPT 3.5 and ML respectively. The chord visualizations summarize the biases that explain the (in)consistencies of (D) human-AI choice and (E) voting input method.
@@
The consistency among the voting input methods is explained by preference towards time discounting factors [41] in participatory budgeting, such as sport and welfare projects. The average relative importance is 25.6% and 17.5% respectively (all with p<0.03), with 48.53%, and 37.5% of higher consistency among supporters in the actual voting (p<0.044). Affect heuristics such as preference for projects that benefit families with children also strongly contributes to the consistency with 58.9% higher consistency among supporters in the actual voting (p<0.032). Figures, LABEL:fig:merged3aLABEL:fig:explaination-top-choice, and Table LABEL:table:error illustrate additional insights about how personal human traits explain the AI top choice and the human-AI consistency of the individual choices.